# SloPalSpeech: A 2,800-Hour Slovak Speech Corpus from Parliamentary Data

### v1

@@ Abstract
Automatic Speech Recognition (ASR) for low-resource languages like Slovak is hindered by the scarcity of training data. To address this, we introduce SloPalSpeech, a new, large-scale Slovak ASR dataset containing 2,806 hours of speech from parliamentary proceedings. We developed a robust processing pipeline to align and segment long-form recordings into clean, 30-second audio-transcript pairs suitable for model training. We use this dataset to fine-tune several OpenAI Whisper models (small, medium, large-v3, and large-v3-turbo), achieving significant Word Error Rate (WER) reductions on standard Slovak benchmarks like Common Voice and FLEURS. For instance, the fine-tuned Whisper-small model’s WER dropped by up to 70%, approaching the baseline performance of the much larger Whisper-large-v3 model. To foster future research in low-resource speech recognition, we publicly release the complete SloPalSpeech dataset, the fully segmented transcripts (60 million words), and all our fine-tuned models.
Keywords: Slovak ASR, Speech recognition, Parliamentary data, Whisper fine-tuning
@@
SloPalSpeech: A 2,800-Hour Slovak Speech Corpus from Parliamentary Data
@@ 1.  Introduction
Automatic Speech Recognition (ASR) has improved significantly with large multilingual datasets and models such as Whisper Radford et al. (2023). However, support for Slovak remains limited. To address this gap, we explore publicly available parliamentary proceedings as a large-scale source of Slovak speech data. We collect and process thousands of hours of recordings and transcripts to create an ASR dataset and fine-tune multiple Whisper models.
Public ASR benchmarks, such as the Hugging Face Open ASR Leaderboard Srivastav et al. (2023), reveal substantial variation among top-performing models from diverse contributors. However, support for low-resource languages like Slovak remains limited. Many leading models either lack Slovak support entirely or are proprietary, restricting their usability for research. From the models shown in Table 1, the first openly available model with Slovak support is CrisperWhisper Zusag et al. (2024), a modified version of OpenAI’s Whisper large-v3.
@@
nvidia/canary-gwen-2.5b 5.63
@@
Table 1: Average WER (%) of top-rated models in Open ASR leaderboard at the time of writing
@@
Models like OpenAI’s Whisper Radford et al. (2023) rely on large, diverse datasets, yet public resources remain scarce for many languages, leading to weaker performance. As shown in Table 2 and Table 3, Whisper achieves much lower WERs on high-resource languages (e.g., English, Russian, German; thousands of training hours) than on low-resource ones (e.g., Slovak, Croatian, Lithuanian; $\sim$100 training hours). Large models can reach 3–5% WER in high-resource settings, whereas low-resource languages often struggle to go below 10%, underscoring the need for more Slovak training data to close this gap.
Model Slovak Croatian Lithuanian
Whisper tiny 77.2 79.0 98.5
Whisper small 33.3 33.4 65.6
Whisper medium 17.3 19.3 41.1
Whisper large‑v2 11.7 13.4 28.1
Hours 90 91 67
Table 2: Whisper WERs on FLEURS on selected low-resource languages
Model English Russian German
Whisper tiny 12.4 31.1 27.8
Whisper small 6.1 11.4 10.2
Whisper medium 4.4 7.2 6.5
Whisper large‑v2 4.2 5.6 4.5
Hours 438k 9.8k 13.3k
@@
Table 3: Whisper WERs on FLEURS on selected high-resource languages
@@ 2.  Objectives
The objective of this work is to address the challenges faced by low-resource languages, as outlined in Section 1, with a particular focus on the Slovak language. Our primary goal is to explore existing resources of Slovak parliamentary speech and construct a large-scale dataset suitable for training speech recognition models. This dataset will consist of audio clips, each no longer than 30 seconds, paired with corresponding textual transcriptions.
Publicly available resources rarely exist in this format, making it likely that preprocessing will be required. As a result, some data may be lost during this process. Therefore, we also aim to investigate the possibility of obtaining fully transcribed textual data, even when no matching audio is available, to maximize the utility of available resources.
The 30-second clip length is chosen in alignment with the Whisper family of models, which we intend to fine-tune using the curated dataset. Our goal is to develop a set of models optimized for Slovak speech recognition. These will include fine-tuned versions of models from Whisper-small up to Whisper-large-v3 (excluding versions v1 and v2). By training models across different sizes, we aim to assess the impact of our dataset on model performance and language-specific accuracy.
@@ 3.  Dataset
Several open-source datasets exist for Slovak, such as CommonVoice Ardila et al. (2019), FLEURS Conneau et al. (2023), and VoxPopuli Wang et al. (2021), but together they offer only about 100 hours of data which is far below the scale required for high-performance ASR. We focus on Slovak parliamentary hearings as a promising resource to bridge this gap, potentially providing thousands of hours of training data and far surpassing existing datasets like those mentioned. These hearings are fully public and regularly transcribed, making them a strong candidate for ASR dataset creation. Audio recordings and their corresponding transcripts are publicly available on various government websites, with coverage dating back to the establishment of the Slovak Republic. This long-term availability presents a valuable opportunity to construct a dataset that qualifies as large-scale. To create this, we must first assess their original availability.
@@ 3.2.  The Joint Czech and Slovak Digital Parliamentary Library
This initiative, established in 2002 in Slovakia, aims to digitize parliamentary documents. Since verbatim transcripts of parliamentary hearings fall within this category, these documents can be easily filtered and accessed in digital form through the portal. The platform features a tree-view interface, which offers an intuitive way to filter transcripts for each session, as illustrated in Figure 2. Under each session, multiple links are provided - each corresponding to the verbatim transcript of a specific meeting day, available in DOCX format.
@@
Figure 2: Example layout from the Digital parliamentary library
@@
We collected all available audio recordings from MediaPortál. Since the frontend download functionality was not working, we extracted the URLs of the HLS streams and used FFmpeg FFmpeg Developers (2025) to download the recordings. Due to storage limitations, we opted to extract only the audio in MP3 format. This resulted in a total of 4512.5 hours of audio.
For the transcripts, we directly downloaded the available DOCX documents. Some documents had to be discarded because the webpage redirected to an HTML page instead of a valid DOCX file. In total, we successfully collected 2111 transcript files.
@@ 3.4.  Parsing
To filter out all non-verbatim content, we needed to segment the transcript based on speaker annotations. However, due to inconsistencies caused by human error (missing commas, inconsistent formatting of names) or edge cases like three-part names, we were unable to rely on a simple split based on commas or regular expressions. Therefore, we adopted a heuristic-based approach using the following conditions:
@@
To support this logic, we scraped a complete list of National Council members from all electoral terms using the official parliamentary website NRSR . From this data, we compiled a comprehensive set of known names, including both first and last names for each member. Each time a bold line was encountered, we checked it against this set. If the required number of name matches was found, the line was classified as a speaker annotation. We also applied a word limit, as we observed that this conditions alone matched some text that was not a speaker annotation but a spoken text. After applying these steps, we achieved a structured format as shown in Listing 4.
Additionally, we needed to filter out all text enclosed in parentheses, as such content typically represents transcriber’s notes. Specifically, we removed all text enclosed in both () and []. To handle cases where parentheses were not properly closed, we implemented a pattern that treated a bracket as the start of the note and a period as the end.
"speaker": "Laššáková, Jana, podpredsedníčka NR SR",
"transcript": "Príjemné, dobré ráno, vážené panie poslankyne, páni poslanci…"
"speaker": kripek, Branislav, poslanec NR SR",
"transcript": akujem pekne, pani predsedajúca. Áno, preto som sa chcel vydýchať…"
Figure 4: JSON example of parsed transcript
@@ 4.1.  Reference Transcript
We first generated a reference transcript with word-level timestamps using the WhisperX framework, which combines Whisper’s transcription capabilities with external alignment models Bain et al. (2023). While Whisper provides reasonably accurate results (with a WER of 12.9% for the largest model, as shown in Table 4), the output still contained typical inaccuracies. However, since it outputs transcripts in 30-second chunks, it enabled effective forced alignment afterward without the limitations previously mentioned.
@@ 4.3.  Segment Construction from Anchors
The anchoring method proved reasonably effective, as most sessions contained a sufficient number of matched words. To construct segments suitable for our ASR dataset, we aimed to group transcript and audio pairs that were no longer than 30 seconds.
@@ 5.  Fine-tuning
With the dataset prepared, we proceeded to fine-tune multiple Whisper models to then be able to assess the impacts of parliamentary speech data on Slovak ASR performance. This section outlines our experimental setup and the fine-tuning strategies adopted for different model sizes, including the adaptations required for the largest variant.
@@ 5.1.  General Setup
Although we had over 2,800 hours of training data, its formal and domain-specific nature raised concerns about catastrophic forgetting. The model could adapt well to parliamentary speech but lose generalization capability for broader Slovak. To assess this, we monitored performance not only on a held-out parliamentary test set but also on subsets of FLEURS and Common Voice. For both of these datasets we produced a subset of 300 samples. These additional benchmarks provided insight into how well the model retained general Slovak transcription capabilities across more diverse contexts. The parliamentary test set was of similar size.
@@ 5.2.  Whisper Small, Medium, Turbo
We applied early stopping between the 2nd and 3rd epoch after observing evaluation plateaus on all of the test sets. None of the models showed signs of overfitting or catastrophic forgetting. All achieved significant improvements on the parliamentary test set, and importantly, also generalized well to broader Slovak speech - gradually yielding notable WER reductions on both the FLEURS and Common Voice subsets during training.
@@ 5.3.1.  Multi-GPU setup
We addressed the GPU limitations by using Fully Sharded Data Parallel (FSDP) Zhao et al. (2023). Although Hugging Face Transformers Wolf et al. (2020) officially support FSDP and multi-GPU training via Accelerate Gugger et al. (2022), we faced persistent compatibility issues - different library versions produced inconsistent runtime errors with the same code. Training was eventually stabilized using version 4.52.4, but evaluation consistently failed due to runtime errors. As a workaround, we used three GPUs for training and a fourth GPU to run inference at selected checkpoints. This differed from the single-GPU setup used for smaller models, where training and evaluation could run together in a single process.

### v2

# SloPal: A 60-Million-Word Slovak Parliamentary Corpus with Aligned Speech and Fine-Tuned ASR Models

@@ Abstract
Keywords: Slovak speech recognition, Whisper fine-tuning, Parliamentary speech corpus, Low-resource ASR, Slavic languages
[…]
SloPal: A 60-Million-Word Slovak Parliamentary Corpus with Aligned Speech and Fine-Tuned ASR Models
@@ 1.  Introduction
Automatic Speech Recognition (ASR) has improved significantly with large multilingual datasets and models such as Whisper Radford et al. (2023). However, support for Slovak remains limited. To address this gap, we leverage publicly available parliamentary proceedings as a large-scale source of Slovak language data. We present SloPal, a 60-million-word corpus of speaker-segmented parliamentary transcripts, alongside SloPalSpeech, a derived aligned speech dataset, and fine-tuned ASR models.
Public ASR benchmarks, such as the Hugging Face Open ASR Leaderboard Srivastav et al. (2023), a snippet of which is shown in Table 1, rank models by Word Error Rate (WER), the standard metric measuring the percentage of incorrectly recognized words. On English, where most evaluation is concentrated, top models achieve closely clustered results with minimal variation. In contrast, support for low-resource languages like Slovak is largely absent, leaving their performance unexplored. Moreover, many leading models either lack Slovak support entirely or are proprietary, restricting their usability for research.
@@
nvidia/canary-qwen-2.5b 5.63
@@
Table 1: Average WER (%) on English test sets of top-rated multilingual models in Open ASR leaderboard at the time of writing
@@
Models like OpenAI’s Whisper Radford et al. (2023) rely on large, diverse datasets, yet public resources remain scarce for many languages, leading to weaker performance. As shown in Table 3, Whisper achieves much lower WERs on high-resource languages (e.g., English, Russian, German; thousands of training hours) than on low-resource ones (e.g., Slovak, Croatian, Lithuanian; $\sim$100 training hours). Large models achieve 3–5% WER on high-resource languages but remain above 10% on low-resource ones—a gap driven primarily by data scarcity rather than model capacity.
Recent work has demonstrated that fine-tuning Whisper on domain-specific or language-specific data can substantially improve performance. Liu et al. (2024) conducted systematic experiments comparing various parameter-efficient fine-tuning approaches for Whisper, showing that targeted adaptation can yield significant improvements even with limited computational resources. Timmel et al. (2024) explored fine-tuning strategies for low-resource settings, confirming that even modest amounts of in-domain data can substantially reduce WER when properly leveraged.
$1.$  Parliamentary Speech Corpora for ASR
Parliamentary proceedings have emerged as a valuable source of speech data for ASR across multiple languages, offering long-form recordings with corresponding transcripts, formal register with clear articulation, and substantial volume. Helgadóttir et al. (2017) pioneered this approach with the Icelandic Althingi corpus, collecting 542 hours of parliamentary speech and establishing methodological foundations for alignment and segmentation of long-form recordings. Solberg and Ortiz (2022) developed the Norwegian Parliamentary Speech Corpus (NPSC), comprising 140 hours of speech, demonstrating a 22.9% relative WER reduction. Kulebi et al. (2022) introduced ParlamentParla, a 600-hour corpus of Catalan parliamentary speech, providing insights into handling speaker diarization and long-form alignment challenges. Rekathati (2023) released RixVox, a Swedish parliamentary corpus with over 5,400 hours of speech, representing the largest single-language parliamentary dataset to date. Virkkunen et al. (2023) released the Finnish Parliament ASR corpus with over 3,000 hours of speech. Stankov et al. (2025) created ParCzech4Speech with 2,695 hours of Czech speech, a West Slavic language closely related to Slovak. More recently, Ljubešić et al. (2024) released the ParlaSpeech collection, an ambitious multi-lingual effort encompassing Croatian (3,110 hours), Polish (1,010 hours), and Serbian (896 hours), demonstrating the scalability of parliamentary corpus construction across Slavic languages. On the textual side, the ParlaMint project Erjavec et al. (2023) created uniformly encoded, linguistically annotated parliamentary corpora for 29 European parliaments, though notably without Slovak or aligned audio. Table 2 summarizes these parliamentary speech corpora alongside SloPalSpeech.
Corpus Language Hours
Althingi Helgadóttir et al. (2017) Icelandic 542
NPSC Solberg and Ortiz (2022) Norwegian 140
ParlamentParla Kulebi et al. (2022) Catalan 600
RixVox Rekathati (2023) Swedish 5,493
Finnish Parl. Virkkunen et al. (2023) Finnish 3,130
ParCzech4Speech Stankov et al. (2025) Czech 2,695
ParlaSpeech-HR Ljubešić et al. (2024) Croatian 3,110
ParlaSpeech-PL Ljubešić et al. (2024) Polish 1,010
ParlaSpeech-SR Ljubešić et al. (2024) Serbian 896
EuroSpeech Pfisterer et al. (2025) 22 languages 50,500
   EuroSpeech-Slovak subset Slovak 2,554
SloPalSpeech (ours) Slovak 2,806
Table 2: Comparison of parliamentary speech corpora for ASR. EuroSpeech is a multilingual corpus with 22 languages totaling 50,500 hours (CER << 20%), including 2,554 hours for Slovak.
At 2,806 hours, SloPalSpeech is the first dedicated Slovak parliamentary speech resource, comparable in scale to ParCzech4Speech (2,695 hours for Czech, a closely related West Slavic language) and the Slovak subset of the concurrent EuroSpeech corpus (2,554 hours). The broader SloPal corpus additionally provides the complete textual record—330,000 speaker-segmented transcripts—making it the largest Slovak parliamentary language resource in any modality.
1.2.  Relation to EuroSpeech
Concurrent with our work, Pfisterer et al. (2025) introduced EuroSpeech, a multilingual parliamentary corpus including 2,554 hours of Slovak speech using a two-stage dynamic alignment achieving $\sim$89% coverage. While EuroSpeech excels at audio alignment across 22 languages, SloPal provides complementary strengths: (1) comprehensive textual coverage [^1] By textual coverage we mean that SloPal includes the full text of all successfully retrieved and parsed transcripts, independent of whether the corresponding audio could be aligned. Some transcript files were unavailable due to download errors (see Section 3.3), and audio recordings cover only forenoon sessions (see Section 3.5). (66M words in 330,000 speaker-segmented transcripts vs. 11,700 segments in EuroSpeech’s publicly available Slovak subset), (2) rich speaker metadata enabling discourse analysis, (3) 30-second segment lengths optimized for Whisper’s architecture, and (4) Slovak-specific fine-tuned models. Our dual-resource design provides complete text coverage for NLP applications while offering 2,806 hours of aligned audio for ASR training. Both resources draw from the same Slovak parliamentary sources (MediaPortál), serving distinct but complementary research needs: EuroSpeech for multilingual ASR benchmarking, SloPal for comprehensive Slovak language and discourse research. We note that Slovak parliamentary audio is publicly available and portions may also appear in other large-scale datasets such as VoxPopuli Wang et al. (2021), which includes European Parliament recordings; SloPal is the first resource to systematically cover the full Slovak National Council proceedings with speaker-level segmentation.
Model Low-resource High-resource
Slovak Croatian Lithuanian English Russian German
Whisper tiny 77.2 79.0 98.5 12.4 31.1 27.8
Whisper small 33.3 33.4 65.6 6.1 11.4 10.2
Whisper medium 17.3 19.3 41.1 4.4 7.2 6.5
Whisper large‑v2 11.7 13.4 28.1 4.2 5.6 4.5
Hours 90 91 67 438k 9.8k 13.3k
@@ 1.2.  Relation to EuroSpeech
Table 3: Whisper WERs (%) on FLEURS for low-resource vs. high-resource languages, illustrating the gap driven by training data availability
@@ 2.  Objectives
The objective of this work is to address the challenges faced by low-resource languages, with particular focus on Slovak. Our primary goal is to create comprehensive resources from parliamentary proceedings: (1) SloPal, a complete corpus of speaker-segmented transcripts capturing the full textual record, and (2) SloPalSpeech, an aligned audio-transcript dataset suitable for ASR model training with segments up to 30 seconds.
Since publicly available resources rarely exist in aligned format, preprocessing inevitably loses some data during alignment. We address this by releasing both the complete textual corpus (SloPal, 60 million words) and the successfully aligned subset (SloPalSpeech, 2,806 hours), maximizing resource utility.
The 30-second segment length aligns with the Whisper family of models, which we fine-tune using SloPalSpeech. Our goal is to develop Slovak-optimized ASR models spanning sizes from Whisper-small to Whisper-large-v3, assessing the impact of parliamentary data on model performance and establishing benchmarks for Slovak ASR.
@@ 3.  Dataset
Several open-source datasets exist for Slovak, such as Common Voice Ardila et al. (2019), a crowd-sourced multilingual speech corpus, FLEURS Conneau et al. (2023), a multilingual speech benchmark derived from translated Wikipedia sentences, and VoxPopuli Wang et al. (2021), a large-scale multilingual speech dataset based on European Parliament recordings. However, together they offer only about 100 hours of Slovak data, which is far below the scale typically required for high-performance ASR.
We focus on Slovak parliamentary hearings to bridge this gap, a source that yields thousands of hours of training data, surpassing existing Slovak datasets by an order of magnitude. These hearings are fully public and regularly transcribed, making them a strong candidate for ASR dataset creation. Audio recordings and their corresponding transcripts are publicly available on various government websites, with coverage dating back to the establishment of the Slovak Republic. This long-term availability presents a valuable opportunity to construct a dataset that qualifies as large-scale.
@@ 3.2.  The Joint Czech and Slovak Digital Parliamentary Library
This initiative, established in 2002 in Slovakia, aims to digitize parliamentary documents. Since verbatim transcripts of parliamentary hearings fall within this category, these documents can be easily filtered and accessed in digital form through the portal. The platform features a tree-view interface, which offers an intuitive way to filter transcripts for each session. Under each session, multiple links are provided - each corresponding to the verbatim transcript of a specific meeting day, available in DOCX format.
@@
We collected all available audio recordings from MediaPortál. Since the frontend download functionality was not working, we extracted the URLs of the HLS (HTTP Live Streaming) streams and used FFmpeg FFmpeg Developers (2025) to download the recordings. Due to storage limitations, we opted to extract only the audio in MP3 format. This resulted in a total of 4096 hours of audio.
For the transcripts, we directly downloaded the available DOCX documents. Some documents had to be discarded because the webpage redirected to an HTML page instead of a valid DOCX file. In total, we successfully collected 2199 transcript files.
@@ 3.4.  Parsing
To filter out all non-verbatim content, we needed to segment the transcripts based on speaker annotations. However, due to inconsistencies introduced during manual transcription (e.g., missing commas, inconsistent formatting of names) and edge cases such as multi-part names, simple rule-based splitting using punctuation or regular expressions proved unreliable. We therefore adopted a heuristic-based approach for identifying speaker annotations using the following conditions:
@@
These thresholds reflect the structure of Slovak parliamentary transcripts: speaker annotations typically contain a surname, first name, and role designation (hence one to three name tokens), and never exceed a short line (hence the 15-word limit), while longer bold passages invariably correspond to section headers or emphasized speech.
To support this logic, we scraped a complete list of National Council members from all electoral terms using the official parliamentary website NRSR . From this data, we compiled a comprehensive set of known names, including both first and last names for each member. Each time a bold line was encountered, we checked it against this set. If the required number of name matches was found, the line was classified as a speaker annotation. During validation, we observed a small number of mismatches caused by missing names in the scraped list, primarily names of Slovak presidents and government ministers who address parliament but are not members. These were added manually to ensure correct speaker identification. Other occasional non-MP speakers (e.g., invited experts) appear rarely in the transcripts and are handled by the same heuristics, since their annotations follow the same bold-text format with name and role.
These heuristics were developed iteratively through manual inspection of the parsing outputs. During development, we repeatedly sampled parsed transcripts and adjusted the conditions until the resulting segmentation produced consistent speaker annotations without introducing obvious errors. In particular, we introduced the word limit after observing that some bold segments matched the name-based conditions but corresponded to emphasized spoken text rather than speaker identifiers.
Additionally, we filtered out all text enclosed in parentheses, as such segments typically represent transcriber’s notes (e.g., descriptions of non-verbal events). Specifically, we removed all text enclosed in both () and []. To handle cases where parentheses were not properly closed, we implemented a fallback rule that treats a bracket as the start of a note and the next period as its end. After applying these steps, the transcripts were transformed into a semi-structured format consisting of alternating speaker annotations and corresponding speech segments.
@@ 4.1.  Reference Transcript
We first generated a reference transcript with word-level timestamps using the WhisperX framework, which combines Whisper’s transcription capabilities with external alignment models Bain et al. (2023). While Whisper provides accurate results (with a WER of 12.9% for the largest model, as shown in Table 4), the output still contained typical inaccuracies. However, since it outputs transcripts in 30-second chunks, it enabled effective forced alignment afterward without the limitations previously mentioned.
@@ 4.2.  Anchor-Based Timestamping
Importantly, this anchor-based approach is language-agnostic: it requires only a Whisper model capable of producing approximate transcripts in the target language and a ground-truth transcript. Because Whisper supports over 90 languages, the method can be applied to parliamentary corpora in any of those languages without training a dedicated forced-alignment model.
@@ 4.3.  Segment Construction from Anchors
The anchoring method was effective: most sessions contained a sufficient number of matched words. To construct segments suitable for our ASR dataset, we aimed to group transcript and audio pairs that were no longer than 30 seconds.
@@ 5.  Fine-tuning
With the dataset prepared, we fine-tuned multiple Whisper models to measure the impact of parliamentary speech data on Slovak ASR performance. This section outlines our experimental setup and the fine-tuning strategies adopted for different model sizes, including the adaptations required for the largest variant.
@@ 5.1.  General Setup
Although we had over 2,800 hours of training data, its formal and domain-specific nature raised concerns about catastrophic forgetting. To assess generalization, we monitored performance during training on held-out evaluation sets: a parliamentary set ( 300 samples) and 300-sample subsets from FLEURS and Common Voice. We selected subsets of 300 samples to keep evaluation computationally feasible without significantly slowing down training. These subsets were sampled from the official test splits; however, they were used solely for early stopping decisions at epoch-level granularity and never for gradient updates. Final benchmarks in Section 6 were conducted on the complete test sets.
@@ 5.2.  Whisper Small, Medium, Turbo
We applied early stopping between the 2nd and 3rd epoch after observing evaluation plateaus. None of the models showed signs of overfitting or catastrophic forgetting. All achieved substantial improvements on the parliamentary evaluation set and importantly also generalized well to broader Slovak speech, yielding notable WER reductions on both FLEURS and Common Voice during training.
@@ 5.3.1.  Multi-GPU setup
We addressed the GPU limitations by using Fully Sharded Data Parallel (FSDP) Zhao et al. (2023). Although Hugging Face Transformers Wolf et al. (2020) officially support FSDP and multi-GPU training via Accelerate Gugger et al. (2022), we faced persistent compatibility issues - different library versions produced inconsistent runtime errors with the same code. Training was eventually stabilized, but evaluation consistently failed due to runtime errors. [^4] Stable training required Transformers v4.52.4; evaluation required a separate process due to FSDP-related runtime errors. As a workaround, we used three GPUs for training and a fourth GPU to run inference at selected checkpoints. This differed from the single-GPU setup used for smaller models, where training and evaluation could run together in a single process.